Please help me with this problem
I do not understand what pi is.Thanks for your help so much.
The circular surface of a table has a diameter of 4 ft. What will cost to have the top refinished if the company charges $6 per square foot for the refinishing? Use 3.14 for piand round your answer to the nearest cent.

Answer by checkley71(8403)

A DIAMETER OF 4 FEET IS TRHE SAME AS A RADIUS OF 2 FEET
AREA=PIR^2
AREA=3.14*R^2
AREA=3.14*4
AREA=12.56 FEET^2.
6*12.56=$75.56 IS THE COST TO REFINISH THE TABLE.